Given
G(t) = 5 − 3t,
write
G(−3 + h) − G(−3)
in simplest form.

Answer by math_helper(2461)


G(-3+h) tells you to substitute "-3+h" for "t" in G(t):
G(-3+h) = 5-3(-3+h) = 5+9-3h = 14-3h

Next find G(-3) by substituting -3 for "t" in G(t)

Take that result and subtract it from G(-3+h) (= 14-3h)

I got an answer of -3h